A gaseous mixture contains 450.0 Torr H2(g), 372.5 Torr N2(g), and 88.1 Torr Ar(g). Calculate the mole fraction, χ, of each of these gases.
Answer #1

The mole fraction of each of the given gases in the gaseous mixture can be calculated as

Xi = P​​​​​​i / P​​​​total ...........(1)

​​​​P​​i = pressure of individual gases (Given)

P​​​​​total = total pressure of the gaseous mixture

Therefore

P​​​​​​total= P​​​​​​N2 + P​​​​​​H2 + P​​​​​​Ar

= (372.5 torr) + ( 450.0 torr) + (88.1 torr)

= 910.6 torr

Therefore, using equation (1) we get,

X​​​​​​H2 = 450.0/ 910.6

= 0.49

X​​​​​​N2​​​​​​​​ = 372.5/ 910.6

= 0.41

X​​​​​​Ar = 88.1/ 910.6

= 0.09